ela is twice as old as her friend lilia. lilia is 7 years older than raquel. ten years ago, ela was three times as old as raquel. how old is lilia now ?
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

ela is twice as old as her friend lilia. lilia is 7 years older than raquel. ten years ago, ela was three times as old as raquel. how old is lilia now ?
------
Equations:
E = 2L
L = R + 7
E-10 = 3(R-10)
---
Substitute for E and R and solve for L:
2L-10 = 3(L-7-10)
2L-10 = 3L-51
L = 41
---
E = 2L = 82
---
R = L-7 = 34
================
Cheers,
stan H.

Answer by Edwin McCravy(20086)

ela is twice as old as her friend lilia. lilia is 7 years older than raquel. ten years ago, ela was three times as old as raquel. how old is lilia now ?
ela is twice as old as her friend lilia.
E = 2L

Lilia is 7 years older than raquel.
L = R + 7

ten years ago, ela was three times as old as raquel [WAS THEN].
E - 10 = 3(R - 10)

So you have the system of three equations in three unknowns:

ìE = 2L
íL = R + 7
îE - 10 = 3(R - 10)


Can you solve that system of equations?  If not, post again 
asking how.

Solution:  Ella is 82, Lelia is 41 and Raquel is 34.


Checking:

Ela, who is 84, is twice as old as her friend Lilia, who is 41. That checks.

Lilia, who is 41, is 7 years older than Raquel, who is 34.  That checks. 

Ten years ago, Ela was 72, which is three times as old as Raquel was then,
for Raquel was 24 then, and 3×24 = 72.  So that checks so we have the 
correct solution.

Edwin